# Explaining Context Length Scaling and Bounds for Language Models

### v2

@@ Appendix D More experiments of LLaMa on another dataset
According to the technical report of LLaMa 3(Grattafiori et al., 2024), the text corpora with number of ‘dirty words’ beyond certain threshold would be filtered out, as proposed in (Raffel et al., 2023). We collect some text corpora online which include forbidden words defined in (Raffel et al., 2023), as text corpora unseen by LLaMa 3. By conducting experiments on it we obtain results similar to Openwebtext subset.
@@ F.1.1 Optimal Context Length Experiments
We use aanogpt(Karpathy, 2022) and train a model with GPT-2(Radford et al., 2019) architecture on a subset of OpenwebText dataset. Our model is the same with GPT-2-124M ($12$-head transformers, $768$-dim feature vector) except that it uses half the transformers layer size ($12\rightarrow 6$) to reduce GPU memory for long contexts. For training, we use the AdamW(Loshchilov & Hutter, 2019) optimizer, an equivalent batch size of $480$ learning rate of $46e-$ 4, weight decay of $11e-$ 1, $2000$ warm-up iterations and $600000$ lr-decaying iterations. We train until overfitting, that is, we validate the validation loss of our model on the validation set every $45$ steps, and choose the best iteration as the min validation loss measured for a single run.
We train the model on a subset of OpenWebText. To be specific, we first select text corpora with context length beyond specific limits larger than the maximum training context length from OpenWebText, then split into Training set and Validation Set. The training set we used to train the models have $12M,25M,50M$ tokens respectively, and the validation set has $134M$ tokens.
Each point in the corresponding figure in Figure 1 and Figure 5 is produced with $8$ AMD MI-250X GPUs (which are similar in performance to Nvidia A100 gpus) trained for $2$ days.
@@ F.1.2 Intrinsic Dimension Experiments
We select long enough text corpora from the Openwebtext dataset. Then, following previous practice(Cheng et al., 2023), we conduct experiments with LLaMa-3.1-8b on $10000$ samples of this subset. We extract the feature representation of the last token in the last layer, as the Intrinsic Representation of samples.
@@ F.2 Synthetic Dataset
We use a large enough MLP on large enough datasets. To be specific, our MLP has four linear layers and two leaky-relu activation layers. The input has shape $context\_length+taskbits\_length$, and output has shape $1$. The hidden layer is of dimension $400$ and middle layer is of dimension $200$. During training, we use the Adam(Kingma & Ba, 2017) Optimizer, batch size is $10000$, learning rate is set to $31e-$ 3, weight decay is $41e-$ 4 and we train until validation loss increases for $25$ epochs, with a maximum epoch number of $200$.
@@
Similar training settings are used for training Synthetic Data for Appendix C.
The experiments for Synthetic Data take around $10$ hours on an Nvidia RTX-3090 gpu.

### v3

@@ Appendix D More experiments of LLaMa on another dataset
According to the technical report of LLaMa 3.1[21], the text corpora with number of ‘dirty words’ beyond certain threshold would be filtered out, as proposed in [23]. We collect some text corpora online which include forbidden words defined in [23], as text corpora unseen by LLaMa 3.1. By conducting experiments on it we obtain results similar to Openwebtext subset.
@@ F.1.1 Optimal Context Length Experiments
We use nanogpt[24] and train a model with GPT-2[25] architecture on a subset of OpenwebText dataset. Our model is the same with GPT-2-124M ($12$-head transformers, $768$-dim feature vector) except that it uses half the transformers layer size ($12\rightarrow 6$) to reduce GPU memory for long contexts. For training, we use the AdamW[26] optimizer, learning rate of $46e-$ 4, weight decay of $11e-$ 1, $1000$ warm-up iterations. For given token number, all models with different context length are trained with same number of iterations, where iteration number equals roughly to $token\_number/(0.1M)$.
We train the model on a subset of OpenWebText. To be specific, we first select text corpora with context length beyond specific limits larger than the maximum training context length from OpenWebText, then split into Training set and Validation Set. The training set we used to train the models have $200M,250M,300M,350M,500M,750M$ tokens respectively, and the validation set has $134M$ tokens.
Experiments presented in Figure 1 and Figure 13 took around $300$ gpu hours on $8$ AMD MI-250X GPUs (which are similar in performance to Nvidia A100 gpus).
@@ F.1.2 Intrinsic Dimension Experiments
We select long enough text corpora from the Openwebtext dataset. Then, following previous practice[12], we conduct experiments with LLaMa-3.1-8b on $10000$ samples of this subset. We extract the feature representation of the last token in the last layer, as the Intrinsic Representation of samples.
@@ F.2 Synthetic Dataset
We use a large enough MLP on large enough datasets. To be specific, our MLP has four linear layers and two leaky-relu activation layers. The input has shape $context\_length+taskbits\_length$, and output has shape $1$. The hidden layer is of dimension $400$ and middle layer is of dimension $200$. During training, we use the Adam[27] Optimizer, batch size is $10000$, learning rate is set to $31e-$ 3, weight decay is $41e-$ 4 and we train until validation loss increases for $25$ epochs, with a maximum epoch number of $200$.
@@
Similar training settings are used for training Synthetic Data for Appendix B.
The experiments for Synthetic Data take around $20$ hours on an Nvidia RTX-3090 gpu.
Appendix G Related Work
G.1 Enlarging Context Length for LMs
Previous work has made attempts to enlarge the context length of Language Models. Work represented by RoPE[1] uses rotary positional embedding to support generalizing LMs to longer context in inference compared to the training process. These work uses modified positional embeddings to model the relative position dependency in attention mechanism.
There is also work about enhancing long context understanding and exploring Scaling Laws for context length[8]. These work utilize an adjusted pretraining and instruction-finetuning process with more long-context data to enhance the models’ ability on long contexts.
Other work modifying architectures has also been proposed to enhance long context modeling or to simplify deployment of long context LLMs. For example, [28] proposes a training-free RazorAttention algorithm to largely compress the KV cache while maintaining performance unchanged.
Architectures and inference methods have been proposed to reduce inference time and memory cost for Language Models, represented by a series of linear transformer or RNN-based methods.[2, 3, 5]. These methods, largely reducing the computational cost and memory usage for long input contexts, have displayed margin ahead of traditional attention-based langauge models for long context inference.
Currently a common practice to train very large Language Models supporting long context is to use pretrain the model with shorter contexts, then finetune them with longer contexts, as presented in tech reports of LLaMa-3[21] and DeepSeek-v3[19].
G.2 Irrelevant Long Context hurts performance of LMs
Besides context length scaling with relevant contexts, previous researches have studied how LLMs perform for long irrelevant contexts. As an example, [7] studies the performance of current LLMs on an adjusted version of ‘needle in a haystack task, where two pieces of key information are embedded into a long text corpora and a question related to both is asked, similar to that presented in Figure 4. The conclusion of these work is that LLMs would perform worse when there is too much irrelevant information.
G.3 Long Context in another field: Time Series Forecasting
Context length, representing the length of input context, is not unique to Nature Language. For time series forecasting, where machine learning plays an important row, there is also work discussing the impact of context length, represented by [9]. These investigations find that there exists an optimal look-back horizon, which increases with dataset size. However, time series datasets are relatively small compared to NLP datasets, and thus whether this conclusion holds on NLP remains an open problem for this work to study.
G.4 Related Theories for Scaling Laws
Since the discovery of Scaling Laws for Large Language Models [14] or even earlier, there has been theoretical work trying to explain why model performance could benefit from more data points and more model parameters. For exmaple, [16] studies the dataset and model scaling from the data manifold perspective.
Specially for Language Models, there is also previous work proposing all kinds of theoretical models. For example, [17] proposes a feature-quant based theory; [13] views the effect of fine-tuning from the intrinsic dimension perspective; [29] proposes to understand scaling with intrinsic dimensions.
We also noticed a preprint version of a concurrent and independent work with ours very recently[30] [^4] Their work is concurrent and independent with our work: the first version of our work is available on Arxiv since Feb.3, 2025, and their first version is available on Arxiv since Mar.6, 2025, thus these two papers are concurrent and independent with each other.. Their work proposed a theoretical framework to explain dependency of successor tokens to previous tokens from bipartite mutual information perspective, and provided a bound of context length scaling laws caused by limited size of hidden states of models (i.e., hidden state with constant size for SSMs might limit long-context performance, while kv-cache which is linear w.r.t. context length for Transformers might not). In our work, we proposed explanations of long context length scaling from Intrinsic Dimension & Entropy perspective, as well as the bound caused by limited dataset (that may also limit Transformers’ performance), which made their work and our work complementary to each other.